OpenStudy (anonymous):

-2x+2/7y=8/7 x being 0

OpenStudy (anonymous):

y=4

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

when x=0 then 2/7y=8/7 2y=8 y=4

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

I know this may sound stupid but how did you get rid of the 7"s

OpenStudy (anonymous):

multiply both side by 7

OpenStudy (anonymous):

dude in which clas r u????????

OpenStudy (anonymous):

7th grade sorry..i don't have anyone to help me..